A reinforced concrete square pole is to withstand a wind shear force of 450.0 kN.

Calculate the relevant cross-sectional side dimensions of the pole for a safety factor

of 9.00. [Ultimate Shear Strength of the reinforced concrete = 125 MPa]


1
Expert's answer
2022-01-21T08:25:32-0500

Given:

"F=450*10^3\:\rm N"

"\sigma_{\max}=125*10^6\:\rm Pa"

"n=9"


The Shear Strength

"\sigma=\frac{F}{A}=\frac{F}{a^2}=n\sigma_{\max}"

The relevant cross-sectional side

"a=\sqrt{F/(n\sigma_{\rm max})}"

"a=\sqrt{450*10^3/(9*125*10^6)}=0.02\:\rm m"